Suppose there is free entry in the market for microphones. The demand for microphones is given by: QD= 2779 -7P. All firms that produce microphones have identical long run average total cost functions given by: ATC = 3528/q + 7 + 8q.

Calculate the long run number of firms in this market.
Answer #1

Answer

The long-run price of a perfectly competitive market is equal to the minimum average total cost.

The output at minimum average total cost is found by differentiating ATC and equating to zero.

dATC/dQ=-3528/q^2+8=0

q^2=3528/8

q^2=441

q=21

P=ATC=3528/21+7+8*21

P=343

---------

the market quantity is

Q=2779-7*343

Q=378

-------

Number of firms =Q/q=378/21=18

the number of firms in the long run is 18